Angle A and angle B are complementary, that is their measurements add up to 90 degrees. Angle B measures 32 degrees more than angle A. What are the measurements of the two angles?
Answer by Cromlix(4381)

Angle A = x
Angle B = x + 32
x + x + 32 = 90
2x + 32 = 90
2x = 90 - 32
2x = 58
x = 29
Angle A = 29 degrees
Angle B = 61 degrees.
Hope this helps.
:-)